OpenStudy (anonymous):

A landscape contractor planted 37 tulip bulbs, 53 daffodil bulbs, and 82 crocus bulbs in each field. He planted bulbs in 174 fields. What is the total number of bulbs the landscape contractor planted? A. 172 B. 346 C. 28,188 D. 29,928

OpenStudy (texaschic101):

37 + 53 + 82 = 172 in each field There are 174 fields... 172 * 174 = 29,928 (D)